1.     Assume a closed economy and no government. Also assume consumption

               C=50-0.8Yd and investment I=80

a)     find the equilibrium output

b)     what is the implied multiplier


1
Expert's answer
2021-08-16T04:36:02-0400

Calculating the equilibrium output


"Y = C +I"

"Y = 50- 0.8Y+ 80"

"(1+0.8)Y = 130"

"1.8Y = 130"

"Y = \frac{130}{1.8} = 72.222"

"Y = 72.222"


Calculating the multiplier

Multiplier "= \frac{1}{1- MPC}"

"MPC = 0.8"

Multiplier = "\frac{1}{0.2}= 5"